Integrated Nanosystems Research Facility

 

Here is just a quick rundown:

STS and E-beam Evap are up.

SEM, Trion, Sputtering Systems, MA6 are down.

I will try and get more details ASAP.

Thanks,
INRF Staff

IMPORTANT: Our Rates for the remainder of this fiscal year have been updated and will be effective 11/01/12.

Please check out the details at:
https://www.inrf.uci.edu/wordpress/wp-content/uploads/inrf-cleanroom-usage-rates.pdf

Thanks,
INRF Staff

Please come to todays meeting in rm 3311 to get the latest equipment status. We would like to appologize that there haven’t been many equipment blog updates and we are working to correct that.

Thanks,
INRF Staff

Today’s 1:30pm user meeting has been cancelled. The next meeting will take place next Wednesday (10/31/12) as usual.

Thanks,
INRF Staff

  • As a reminder the facility will experience a power shut down by Facilities on Oct 6th from 8am – 5pm. This will impact all the equipment within the INRF facility. The equipment will have to be started up again and we will do our best to have it up by Monday.
  • The Chill water upgrade that supports the thermal evaporator, STS, Plasma therm – the plumbing and filter reconfiguration has been completed.
  • Unfortunately the recirculating pump is too small to support all three tools at the same time and will have to be replaced. At the present time the Plasma therm and Thermal evaporator are back online.
  • STS-ICP will remain off line until the new pump is installed on the chill water system. We are also in the process of replacing the low conductivity pump that supports the RF system.
  • ION milling – We performed a test run on the system Wednesday evening and are still experiencing problems with emission current. The problems seem to be located in the power supply. We will continue looking into the problem as time allows.

Thanks,
INRF Staff

Jake will be out of the lab today and Friday (9/27/12 – 9/28/12). He will be not be available until he returns Monday.

Thanks,
INRF Staff

Equipment:

  • E-beam is finished. The staff is using it to catch up on foundry work and it will be available Monday for users. We still need to dissemble and clean the emitter and we plan to purchase a replacement so we have one in stock to minimize down time.
  • Thermal evaporator is back online. Bill Tang’s group will be using it to deposit Ti.
  • We are planning to address the chill water system with the STS starting Monday. This will impact the Thermal Evaporator, STS and Plasma Therm. To revamp the existing chill water it will take five working days.
  • The replacement lamp for the Flood Exposure in the BiON came in yesterday and will be replaced this afternoon. The system should be up and running later today.

Thanks,
INRF Staff

As you might have noticed, we have been taking down some of the coveralls to be cleaned every week. Every Monday one of our staff inspects the coveralls and pulls the ones that need cleaning or coveralls from inactive users. If you find that your suit is missing, please see Ryan (x2819) at the front desk and he will issue you a new one. In the event Ryan is unavailable, Brittany or Alana should be able to issue you a new suit. We hope this clears things up.

 

Thanks,
INRF Staff

The STS is still experiencing high temperature for the dry pump and also a circulation issue with the chiller that supports the RF network. To correct the dry pump water issue we will have to replace the existing copper lines due to build-up. We have no estimated completion date at this time.

Thanks,
INRF Staff

  • The annual maintenance on the DI water systems for both INRF and BiON was completed. The maintenance consisted of treating the system with chlorine and an inhibitor to remove any bacteria. We confirmed the resistivity was back to 17 megohm in INRF but we still need to verify the BiON. The BiON should be useable Friday.
  • The STS is experiencing chill water issues for the dry mechanical pump. The pump tripped out on high temperature. We need to disassemble the chill water line to decontaminate to get operating temperatures back to normal. It should be ready by Monday.
  • E-beam evaporator: The high vac was disassembled, cleaned and O rings replace. High vac valve is opening and closing normally. We just received the chamber shrouding from the cleaning company. We need to finish cleaning the shutter actuator, reassemble, then re-gen the cryo pump. Possibly ready by Tuesday.
  • XeF2 pulse etcher chamber has been cleaned. We need to reconnect the pumping line to the chamber and verify the system is operating within normal parameters. It should be ready by Friday.
  • ION milling: We are verifying interconnections in the gun. We’ve identified one bad connection leading to the chamber feed through connector. The connector needs to be repaired.
  • Oriel flood exposure still working on high voltage power supply.
  • MA56 mercury lamp needs to be replaced and should be ready by Friday.
  • AB&M flood exposure mercury lamp needs to be replaced. We expect it to be back online by Friday.

Thanks,
INRF Staff